INC000100664395 - P3 - Single Exchange - LSERI - Erith - Loss of voice and data services - VM Backau

Details:

Current Status:Closed
We've identified an issue impacting customer connectivity.

Locations impacted:

LSERI - Erith

Service and customer volumes impacted:

Ethernet: 16

MPF/SMPF: 140

Below is a summary of our initial findings and next steps:

A Virgin Media backhaul is down, a fault has been raised to them to investigate.

We understand the impact this may have caused and appreciate your patience and understanding.

Work Log Submit Date : 11/02/2024 21:32:49

Virgin Media have advised they have lost communications to their Kit in the Exchange and have assigned an engineer to investigate.

We'll continue to keep you informed and updated on any further developments.

Work Log Submit Date : 11/02/2024 22:34:44

Virgin Media have advised there is issue in Lewisham area which may have a common cause with the Loss of service on this Exchange.

We'll continue to keep you informed and updated on any further developments.

WorkLog Submit Date : 12/02/2024 00:18:22

Virgin Media have advised they have an engineer on site at Bexley Heath testing fibres.

We'll continue to keep you informed and updated on any further developments.

WorkLog Submit Date : 12/02/2024 00:36:22

03:19 12/02/2024 Virgin Media have advised the engineer on site has been joined by an ERS Splicer to help with the investigation.


We'll continue to keep you informed and updated on any further developments.

Latest from Virgin Media is that their Engineers are still on site working to find the damage fibres. At the same time they are also looking at re-routing services on spare fibres.

We'll continue to keep you informed and updated on any further developments.

WorkLog Submit Date : 12/02/2024 04:29:02

Virgin Media have advised that their engineers, working with ERS Splicers, have now identified good alternate fibres out of VM Deptford, and that spicing is underway but as this is not a like-for-like splice, this work may take some time and they do not have an estimated time of repair (ETR) yet.

We'll continue to keep you informed and updated on any further developments.

Work Log Submit Date : 12/02/2024 06:00:09

Latest Update:

Virgin media have advisedtheir engineers are still working on site with the splicers.

We currently do not have an ETR for the fix but we will keep you updated as soon as we have more information.

WorkLog Submit Date : 12/02/2024 08:46:48


Next Update: 12/02/2024 12:00:00

Latest Update:

Virgin media have advised the fibres are currently being moved over to spare Fibres.

We currently do not have an ETR for the fix but we will keep you updated as soon as we have more information.


WorkLog Submit Date : 12/02/2024 11:51:04


**RESTORATION**

Services have recovered. We're still awaiting confirmation that the issue has been resolved. As soon as we get it we will let you know.

Latest Update:

Services have recovered. We're still awaiting confirmation that the issue has been resolved. As soon as we get it we will let you know.

WorkLog Submit Date : 12/02/2024 14:49:28

Next Update: 12/02/2024 18:00:00

We're happy to report that service has remained stable and we'll now close this incident.

Work Log Submit Date : 12/02/2024 19:58:05

Start date:21:37 11th Feb 2024
End date:20:05 12th Feb 2024